OpenStudy (anonymous):

What is the slope of a line that passes through the point (−1, 1) and is parallel to a line that passes through (4, 6) and (−1, −4)

OpenStudy (zepp):

They only ask for the slope, and two parallel functions has the ______ slope.

OpenStudy (anonymous):

so I should solve the slope of the 4,6 and -1,-4

OpenStudy (zepp):

Yes, that's right. :)

OpenStudy (zepp):

Since two parallel functions have the same slope.

OpenStudy (anonymous):

Slope of line that passes through (4,6) and (-1,-4) is (6-(-4))/(4-(-1))=10/5=2. Since lines are parallel then their slopes are equal. SO, required line has form y=2x+b. To find b use the fact that line passes through (-1,1): 1=2*(-1)+b -> b=3 Answer: y=2x+3

OpenStudy (zepp):

\(\huge m=\frac{y2-y1}{x2-x1}\)

OpenStudy (zepp):

@simamura, the answer is 2 :)

OpenStudy (anonymous):

ok I solved and got 2/-3

OpenStudy (anonymous):

@zepp you're right.

OpenStudy (zepp):

\(\huge m = \frac{-4-6}{-1-4}=\frac{-10}{-5}=2\)